# Bayesian Perspective for Orientation Estimation in Cryo-EM and Cryo-ET

### v1

@@ 2 Problem formulation and the MMSE orientation estimator
Throughout this paper, we use $g$ to denote both the rotation operator and its corresponding matrix representation. For instance, $g$ can be represented by a $3\times 3$ rotation matrix in 3D. The intended meaning will be clear from the context, and this slight abuse of notation should not cause confusion.
@@ 2.1 Mathematical model for orientation estimation
Let $V(x):\mathbb{R}^{n}\to\mathbb{R}$ be a structure that represents the reference projection $(n=2)$ or volume $(n=3)$, and let $\mathsf{G}$ be the group of 2D or 3D rotations correspondingly. For example, when $n=3$, $V$ corresponds to the continuous 3D electron density, as commonly used in cryo-EM or cryo-ET. We emphasize that the methodology can be readily extended to any compact group, which is a subgroup of the orthogonal group in $\mathbb{R}^{n}$; a notable example is the group of circular shifts [9]. Let $\Pi$ be a known linear operator comprising the discretization sampling operator and potentially a tomographic projection operator. We thus model a single measurement $y$ as
$y=\Pi(g^{-1}\cdot V)+\varepsilon\in\mathbb{R}^{d},$ (2.1)
where $d$ is the total dimension of measurement space, $g\sim\Lambda$ is an unknown random element of $\mathsf{G}$ following a possibly non-Haar (i.e., non-uniform) distribution $\Lambda$ over $\mathsf{G}$, and $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{d\times d})$ is the standard Gaussian measurement noise vector with zero mean and variance level of $\sigma^{2}$. For simplicity, we omit $\Pi$ when it only represents the discretization sampling operator.
Modeling structural uncertainty.
In real-world scenarios, particles infrequently conform to an exact, uniform structure. Instead, they often exhibit slight variations due to different deformations or conformational changes (e.g., [55]). To capture this inherent variability, we introduce a more nuanced model for the volume. Rather than assuming a fixed structure, we represent the volume with a mean $\overline{V}$ and incorporate uncertainty through a diagonal covariance matrix $\Sigma=\operatorname{diag}(\tau_{1}^{2},...,\tau_{d}^{2})$. This covariance matrix can be obtained by discretizing the covariance structure of a pre-specified isotropic Gaussian process $W$ in the $d$ dimensional measurement space, where this Gaussian process reflects the user’s assumptions or understanding of structural uncertainty. The isotropy assumption reflects the idea that structural variations are equally likely in all directions. In other words, we assume $V$ has the following flexible structure
$V(x)\mathrm{d}x=\overline{V}(x)\mathrm{d}x+\mathrm{d}W(x),~{}\text{ for }~{}x\in\mathbb{R}^{3},$ (2.2)
where $\overline{V}$ is a given mean function, and $W$ is a given isotropic Gaussian process over $\mathbb{R}^{3}$, representing possible local conformation differences between $V$ and $\overline{V}$. Here, isotropic means that its statistics remain invariant under rotations; that is, for any $g\in\mathsf{G}$, we have $W(gx)=W(x)$. Combining (2.1) and (2.2) yields the simplified model
$y=\Pi(g^{-1}\cdot\overline{V})+\overline{\varepsilon}+\varepsilon\in\mathbb{R}^{d},$ (2.3)
where $\overline{\varepsilon}\sim\mathcal{N}(0,\mathrm{diag}(\tau_{1}^{2},\ldots,\tau_{d}^{2}))$ with a known variance structure.
@@
We present three typical examples of this model. In all cases, the goal is to estimate $g$ given the sample $y$, the structure $\overline{V}$, and the noise levels $\sigma^{2}$ and $\{\tau_{1}^{2},\ldots,\tau_{d}^{2}\}$.
@@
    2D template matching. In this case, $\Pi$ is solely the discretization sampling operator, $d=N\times N$ with $N$ the grid size of 2D projection images, $g\in\mathsf{G}=\mathrm{SO}(2)$ is a 2D in-plane rotation, and $\overline{V}$ is a given 2D template image. Here, $V$ represents a noisy version of $\overline{V}$, potentially corrupted due to averaging or other pre-processing procedures.
@@
    Rotation estimation in cryo-EM. Here, we consider a special case of (2.3) where $\Pi$ comprises both the sampling and tomographic projection operators, $d=N\times N$ with $N$ the grid size of 2D projection images, $g\in\mathsf{G}=\mathsf{SO}(3)$ is a 3D rotation, and $\overline{V}$ is a given 3D volume representing a known reference 3D structure or a well-grounded structure from prior data analyses.
@@
    3D structure alignment in cryo-ET. In this scenario, we consider a special case of (2.3) where $\Pi$ is solely the discretization sampling operator, $d$ corresponds to the total dimension of 3D subtomograms, $g\in\mathsf{G}=\mathsf{SO}(3)$ is a 3D rotation, and $\overline{V}$ is a given 3D volume. Notably, the 3D alignment problem is also a critical step in the computational pipeline of cryo-EM, see, e.g., [50, 17].
@@
As the MMSE estimator is tightly related to given loss functions, we present two candidate metrics over the 3D orthogonal group $\mathsf{SO}(3)$. It is important to note that our framework is flexible and can accommodate other loss functions over any group $\mathsf{G}$. For a more comprehensive discussion on metrics for rotations, we direct readers to the works of [18, 21]. While similar metrics and estimation procedures apply to the $\mathrm{SO}(2)$ estimation problem, we omit detailed discussion for brevity.
@@
    $d_{F}(g_{1},g_{2})\triangleq\|g_{1}-g_{2}\|_{F}=\sqrt{\mathrm{tr}\Big{(}{\left(g_{1}-g_{2}\right)^{\mathsf{T}}\left(g_{1}-g_{2}\right)}\Big{)}},$ (2.4)
    where $\left\|\cdot\right\|_{F}$ represents the matrix Frobenius norm, and $\mathrm{tr}(\cdot)$ is the trace of a matrix. This metric is easy to compute and analyze, however, it does not take into account the group structure of rotations.
@@
    where $\mathrm{tr}(\cdot)$ is the trace of a matrix. This metric reflects the shortest path between $g_{1}$ and $g_{2}$ over the manifold of 3D rotations.
2.2 The MMSE and MAP orientation estimators
Before introducing the specific estimator for our model (2.3), it is instructive to briefly revisit the general Bayesian framework and the concept of the Bayes estimator. This will provide the necessary foundation for understanding the development and analysis of our proposed MMSE estimator, as well as its superior statistical properties compared to the widely used MAP estimator.
Overview of Bayesian framework and the Bayes estimator.
Suppose we aim to estimate some true rotation $g\in\mathsf{G}$ drawn from a known prior distribution $\Lambda$ from data. Let $\hat{g}$ be an estimator of $g$ based on a measurement $y$ and let $\mathcal{L}\left(g,\hat{g}\right)$ be a loss function (for example, the chordal and geodesic distances). The Bayes risk of $\hat{g}$ is defined as $\mathbb{E}_{\Lambda}[\mathcal{L}\left(g,\hat{g}\right)]$, where the expectation is taken over the data generating process of $y$ given $g$ and the prior distribution $\Lambda$ of $g$. The Bayes estimator [27, Chapter 4, Theorem 1.1] is defined as the estimator that minimizes the Bayes risk among all estimators, i.e.,
Equivalently, it is the estimator that minimizes the posterior expected loss $\mathbb{E}_{\Lambda}[\mathcal{L}\left(g,\hat{g}\right)|y]$, where the expectation is taken over the posterior distribution of $g$ given the measurement $y$, and the known volume $\overline{V}$, with the prior $\Lambda$, i.e.,
$\hat{g}_{\mathsf{Bayes}}=\underset{\hat{g}}{\operatorname*{arg\,min}}\ \mathbb{E}_{\Lambda}\left[\mathcal{L}\left(g,\hat{g}\right)|y,\overline{V}\right],$ (2.7)
with
${\mathbb{P}}_{\Lambda}(g|y,\overline{V})\mathrm{d}g=\frac{{\mathbb{P}}(y|g)\mathrm{d}\Lambda(g)}{\int_{\mathsf{G}}{\mathbb{P}}(y|g)\mathrm{d}\Lambda(g)}.$
The posterior distribution of the rotation $g$ given the observation $y$ under model (2.3).
To introduce the MMSE estimator corresponding to the model (2.3), we first compute the posterior distribution of $g$ given $y$ and all the additional parameters $\overline{V}$, $\sigma^{2}$ and $\{\tau_{1}^{2},\ldots,\tau_{d}^{2}\}$. We can rewrite (2.3) as
$y=$ $\Pi(g^{-1}\cdot\overline{V})+\bar{\varepsilon}+\varepsilon=\Pi(g^{-1}\cdot\overline{V})+\tilde{\varepsilon},$ (2.8)
where $\tilde{\varepsilon}:=\bar{\varepsilon}+\varepsilon\sim\mathcal{N}(0,\tilde{\Sigma})$ with $\tilde{\Sigma}=\mathrm{diag}(\tilde{\tau}_{1}^{2},\ldots,\tilde{\tau}_{d}^{2})$ and $\tilde{\tau}_{i}^{2}=\tau_{i}^{2}+\sigma^{2}$. Therefore, we obtain the conditional likelihood density
${\mathbb{P}}(y|g,\overline{V})=\frac{1}{(2\pi)^{d/2}\mathrm{det}(\tilde{\Sigma})^{1/2}}\exp\Big{(}-\frac{1}{2}(y-\Pi(g^{-1}\cdot\overline{V}))^{\mathsf{T}}\tilde{\Sigma}^{-1}(y-\Pi(g^{-1}\cdot\overline{V}))\Big{)}.$
Note that $g$ has the prior distribution $\Lambda$. Applying the Bayes’ law, the posterior distribution of $g$ given $y$ is
${\mathbb{P}}(g|y,\overline{V})\mathrm{d}g=\frac{\exp\Big{(}-\frac{1}{2}(y-\Pi(g^{-1}\cdot\bar{V}))^{\mathsf{T}}\tilde{\Sigma}^{-1}(y-\Pi(g^{-1}\cdot\overline{V}))\Big{)}\mathrm{d}\Lambda(g)}{\int_{\mathsf{G}}\exp\Big{(}-\frac{1}{2}(y-\Pi(g^{-1}\cdot\overline{V}))^{\mathsf{T}}\tilde{\Sigma}^{-1}(y-\Pi(g^{-1}\cdot\overline{V}))\Big{)}\mathrm{d}\Lambda(g)}.$ (2.9)
@@ 2.2 The MMSE and MAP orientation estimators
Recalling the maximum cross-correlation method we mentioned earlier, we now connect it with our mathematical model (2.8) and introduce the corresponding MAP estimator. Assume, in this explicit model with $\mathsf{G}=\mathsf{SO}(3)$, $\Lambda$ is the uniform distribution and all $\tilde{\tau}_{i}$’s are equal to $\tilde{\tau}$. Then, the rotation $g$ that minimizes the distance between the corresponding projected rotated volume $\Pi(g^{-1}\cdot\bar{V})$ and the observation $y$ is exactly the maximum a posteriori (MAP) estimator defined as
$\hat{g}_{\mathsf{MAP}}:=$ $\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}{\mathbb{P}}(g|y,\bar{V})$
$=$ $\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}\exp\Big{(}-\|y-\Pi(g^{-1}\cdot\overline{V})\|^{2}/2\tilde{\tau}^{2}\Big{)}$
$=$ $\operatorname*{arg\,min}_{g\in\mathsf{SO}(3)}\|y-\Pi(g^{-1}\cdot\overline{V})\|^{2},$ (2.10)
where the first equality holds because the denominator in (2.9) does not depend on $g$, and additionally, $\Lambda$ is the uniform distribution and $\tilde{\Sigma}=\tilde{\tau}^{2}I_{d\times d}$ for the numerator. In absence of the tomographic projection (e.g., in cryo-ET), the MAP estimator further simplifies to
$\hat{g}_{\mathsf{MAP}}=\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}y^{\mathsf{T}}(g^{-1}\cdot\overline{V}),$
which corresponds to the rotation $g$ such that the rotated structure maximizes the correlation with $y$. This estimator is also frequently used in cryo-EM, under the assumption that the norm of $\Pi(g^{-1}\cdot\overline{V})$ is approximately constant for all $g$. This estimator can be approximated by performing a search over a pre-defined grid of 3D rotations, selecting the rotation that minimizes the distance between the measurement $y$ and the projected rotated volume $\Pi(g^{-1}\cdot\overline{V})$. This approach forms the basis of standard practices in cryo-EM and cryo-ET.
@@
For any general rotational distribution $\Lambda$, general variance structure of $\tilde{\varepsilon}$, and loss function $\mathcal{L}(\cdot,\cdot)$ over $\mathsf{SO}(3)$, the MAP estimator can be further improved by the corresponding Bayes estimator which minimizes the posterior expected loss $\mathbb{E}_{\Lambda}[\mathcal{L}(g,\cdot)|y,\overline{V}]$. In particular, following the definition (2.7) of the Bayes estimator, and for the chordal distance (2.4), where $\mathcal{L}(\cdot,\cdot)=d_{F}^{2}(\cdot,\cdot)$, the Bayes estimator is equivalent to the MMSE estimator [27, Chapter 4, Corollary 1.2] :
$\hat{g}_{\mathsf{MMSE}}:=$ $\underset{\hat{g}}{\operatorname*{arg\,min}}\ \mathbb{E}_{\Lambda}\left[d_{F}^{2}\left(g,\hat{g}\right)|y,\overline{V}\right]=\mathbb{E}_{\Lambda}[g|y,\overline{V}]$
$=$ $\frac{\int_{\mathsf{SO}(3)}g\exp\Big{(}-\frac{1}{2}(y-\Pi(g^{-1}\cdot\overline{V}))^{\mathsf{T}}\tilde{\Sigma}^{-1}(y-\Pi(g^{-1}\cdot\overline{V}))\Big{)}\mathrm{d}\Lambda(g)}{\int_{\mathsf{SO}(3)}\exp\Big{(}-\frac{1}{2}(y-\Pi(g^{-1}\cdot\overline{V}))^{\mathsf{T}}\tilde{\Sigma}^{-1}(y-\Pi(g^{-1}\cdot\overline{V}))\Big{)}\mathrm{d}\Lambda(g)},$ (2.11)
where the expectation is taken over the posterior density (2.9). Notably, $\hat{g}_{\mathsf{MMSE}}$ may not be a valid rotation operator in general. Thus, we use the orthogonal Procrustes procedure to obtain a valid rotation, as discussed next. We refer to the estimator obtained through this procedure also as $\hat{g}_{\mathsf{MMSE}}$, provided it causes no misunderstanding. We summarize the numerical procedure for calculating the MMSE estimator for rotation estimation in Algorithm 1.
Orthogonal Procrustes.
It is important to note that $\hat{g}_{\mathsf{MMSE}}$, as defined in (2.11), is not necessarily a valid operator in $\mathsf{SO}(3)$. Consequently, to obtain a valid rotation operator, given a matrix $A$, the objective is to find a matrix representation $\Omega\in\mathsf{SO}(3)$ that is closest to matrix $A$ in terms of the Frobenius norm metric by solving
@@
This problem is called the Orthogonal Procrustes in the linear algebra literature and it solved efficiently by singular value decomposition (SVD) of the matrix $A$ [16]. In particular, given the SVD of $A$, represented by $U\Sigma V^{\mathsf{T}}$, the solution can be expressed as $U\Sigma^{\prime}V^{\mathsf{T}}$, where $\Sigma^{\prime}$ is derived from $\Sigma$ by setting the smallest singular value to $\det(UV^{\mathsf{T}})$ and all other singular values to 1, so that the determinant of the solution is guaranteed to be positive as desired.
The MMSE and MAP estimators in the high SNR regime.
The next proposition shows that the MMSE and MAP estimators converge in the high SNR regime (i.e., $\sigma\to 0$), as demonstrated empirically in Figures 3, 4, and 5. This implies that the MMSE estimator’s superior statistical properties are most advantageous in low SNR conditions, which are common in structural biology applications like cryo-EM and cryo-ET [7]. In these low SNR environments, the MMSE estimator consistently outperforms its MAP counterpart. Assume $\tau_{i}^{2}=0$ as part of the model introduced in (2.3). Thus, the noise covariance matrix is given by $\Sigma=\sigma^{2}I_{d\times d}$. In the following sections, we maintain this setting for simplicity, as it is sufficient to convey our main arguments.

### v2

@@ 2 Problem formulation and the MMSE orientation estimator
Throughout this paper, we use $g$ to denote both the rotation operator and its corresponding matrix representation. For instance, $g$ can be represented by a three by three rotation matrix in 3D. The intended meaning will be clear from the context, and this slight abuse of notation should not cause confusion.
@@ 2.1 Mathematical model for orientation estimation
We consider a unified framework for modeling measurement processes, encompassing problems such as 2D template matching, orientation estimation in cryo-EM, and subtomogram averaging in cryo-ET. To better focus on the core aspects of our methodology, we omit certain physical effects, such as CTF and in-plane shifts, in this formulation. A complete mathematical model of the cryo-EM imaging process, incorporating these effects, is provided in Appendix A.
We begin with a continuous-domain formulation. Let $V(x):\mathbb{R}^{n}\to\mathbb{R}$ denote a reference structure, where $n=2$ for 2D images and $n=3$ for 3D volumes. For example, when $n=3$, $V$ corresponds to the continuous 3D electron density, as commonly used in cryo-EM or cryo-ET. Let $\mathsf{G}\subset\mathrm{SO}(n)$ be a compact group of rotations correspondingly, and let $g\in\mathsf{G}$ be an unknown transformation drawn from a distribution $\Lambda$ over $\mathsf{G}$.
We denote the measurement as a function $y(x^{\prime}):\mathbb{R}^{m}\to\mathbb{R}$, where $m\leq n$ accounts for possible dimension reduction due to projection (e.g., $m=2$ in cryo-EM). Then, the continuous measurement model is given by
$y(x^{\prime})=\Pi\Big((g\circ V)(x)\Big)+\varepsilon(x^{\prime}),\qquad x\in\mathbb{R}^{n}\text{ and }x^{\prime}\in\mathbb{R}^{m},$
where $(g\circ V)(x):=V(g^{-1}x)$ is the rotated structure, $\Pi$ is a known linear operator (such as a tomographic projection or identity), and $\varepsilon$ denotes additive noise in the measurement domain.
In practice, measurements are only available in discretized form due to finite resolution. In the discrete setting, we assume that the operator $\Pi$ also incorporates sampling onto a grid of size $d$, i.e., it includes both projection (if applicable) and discretization. The resulting discrete measurement model becomes
$y_{i}=\Pi(g_{i}\circ V)+\varepsilon_{i},$ (2.1)
where $y_{i}\in\mathbb{R}^{d}$, and $\varepsilon_{i}\sim\mathcal{N}(0,\Sigma)$ is Gaussian noise with $d\times d$ positive-definite covariance matrix $\Sigma$. To simplify notation, we omit $\Pi$ when it solely represents the discretization sampling operator and use the same symbol $y$ to refer to the discretized measurement throughout the paper, unless stated otherwise.
@@
We present three typical examples of this model. In all cases, the goal is to estimate $g$ given the sample $y$, the structure V{V}, and the covaraince matrix $\Sigma$.
@@
    2D template matching. In this case, $\Pi$ is solely the discretization sampling operator, $d=N\times N$ with $N$ the grid size of the 2D image, $g\in\mathsf{G}=\mathrm{SO}(2)$ is a 2D in-plane rotation, and V{V} is a given 2D template image.
@@
    Rotation estimation in cryo-EM. Here, we consider a special case of (2.1) where $\Pi$ comprises both the sampling and tomographic projection operators, $d=N\times N$ with $N$ the grid size of 2D projection images, $g\in\mathsf{G}=\mathsf{SO}(3)$ is a 3D rotation, and V{V} is a given 3D volume representing a known reference 3D structure or a well-grounded structure from prior data analyses.
@@
    3D structure alignment in cryo-ET. In this scenario, we consider a special case of (2.1) where $\Pi$ is solely the discretization sampling operator, $d$ corresponds to the total dimension of 3D subtomograms, $g\in\mathsf{G}=\mathsf{SO}(3)$ is a 3D rotation, and V{V} is a given 3D volume. Notably, the 3D alignment problem is also a critical step in the computational pipeline of cryo-EM, see, e.g., singer2024alignment ; harpaz2023three .
2.2 Preliminaries
Before introducing the specific estimator for the model specified in (2.1), it is instructive to briefly revisit the general Bayesian framework and the concept of the Bayes estimator. This will provide the necessary foundation for understanding the development and analysis of the proposed MMSE estimator, as well as its statistical properties superior to those of the widely used MLE estimator.
Overview of the Bayesian framework and the Bayes estimators.
[…]
Equivalently, it is the estimator that minimizes the posterior expected loss $\mathbb{E}_{\Lambda}[\mathcal{L}\left(g,\hat{g}\right)|y]$, where the expectation is taken over the posterior distribution of $g$ given the measurement $y$, and any other known parameter such as volume V{V}, with the prior $\Lambda$. For the case of $G=\mathsf{SO}(3)$, it is given explicitly by
$\hat{g}_{\mathsf{Bayes}}$ $=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\ \mathbb{E}_{\Lambda}\left[\mathcal{L}\left(g,\hat{g}\right)|y,{V}\right]$
$=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\int_{\mathsf{SO(3)}}\mathcal{L}\left(g,\hat{g}\right){\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g,$ (2.3)
where by Bayes’ law, we have,
${\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g=\frac{{\mathbb{P}}(y|g,{V})\mathrm{d}\Lambda(g)}{\int_{\mathsf{G}}{\mathbb{P}}(y|g,{V})\mathrm{d}\Lambda(g)}.$ (2.4)
The posterior distribution of the rotation $g$ given an observation $y$ under model (2.1).
To introduce the MMSE estimator corresponding to the model (2.1), we first compute the posterior distribution of $g$ given $y$ and all the additional parameters V{V}, and $\Sigma$. We obtain the conditional likelihood density
${\mathbb{P}}(y|g,{V})=\frac{1}{(2\pi)^{d/2}\mathrm{det}({\Sigma})^{1/2}}\exp\Big(-\frac{1}{2}(y-\Pi(g\circ{V}))^{\mathsf{T}}\Sigma^{-1}(y-\Pi(g\circ{V}))\Big).$ (2.5)
Note that $g$ follows the underlying prior distribution $\Lambda$. Applying Bayes’ law, the posterior distribution of $g$ given $y$ is
${\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g=\frac{\exp\Big(-\frac{1}{2}(y-\Pi(g\circ{V}))^{\mathsf{T}}\Sigma^{-1}(y-\Pi(g\circ{V}))\Big)\mathrm{d}\Lambda(g)}{\int_{\mathsf{G}}\exp\Big(-\frac{1}{2}(y-\Pi(g\circ{V}))^{\mathsf{T}}\Sigma^{-1}(y-\Pi(g\circ{V}))\Big)\mathrm{d}\Lambda(g)}.$ (2.6)
@@ 2.2 Preliminaries
As the Bayes estimator is closely related to the given loss function $\mathcal{L}$, we present two candidate metrics on the 3D special orthogonal group $\mathsf{SO}(3)$. Here we represent any rotation $g\in\mathsf{SO}(3)$ in its natural three by three matrix representation.
@@
    $d_{F}(g_{1},g_{2})\triangleq\|g_{1}-g_{2}\|_{\text{F}}=\sqrt{\mathrm{tr}\Big({\left(g_{1}-g_{2}\right)^{\mathsf{T}}\left(g_{1}-g_{2}\right)}\Big)},$ (2.7)
    where $\left\|\cdot\right\|_{\text{F}}$ represents the matrix Frobenius norm, and $\mathrm{tr}(\cdot)$ is the trace of a matrix. This metric is easy to compute and analyze; however, it does not take into account the group structure of rotations.
@@
    where $\mathrm{tr}(\cdot)$ is the trace of a matrix. This metric reflects the shortest path between $g_{1}$ and $g_{2}$ in the 3D rotation manifold.
It is important to note that the framework is flexible and can accommodate other loss functions over any group $\mathsf{G}$. For a more comprehensive discussion of rotation metrics, we direct readers to the work of hartley2013rotation ; huynh2009metrics . Similar metrics and estimation procedures apply to the $\mathrm{SO}(2)$ estimation problem, we omit a detailed discussion for brevity.
2.3 The MLE, MAP and MMSE orientation estimators
In the following presentation, we focus on independent and identically distributed Gaussian noise with $\Sigma=\sigma^{2}I_{d\times d}$. These noise statistics are commonly employed in modern software tools kimanius2021new . While the proposed framework can incorporate more advanced noise models, we restrict attention to this case for clarity of presentation and to simplify the theoretical analysis.
The MLE estimator.
Recalling the maximum cross-correlation method we mentioned earlier, we now connect it with the mathematical model in (2.1) and introduce the corresponding MLE estimator. The rotation $g$ that minimizes the distance between the corresponding projected rotated volume $\Pi(g\circ{V})$ and the observation $y$ is exactly the MLE estimator defined as
$\hat{g}_{\mathsf{MLE}}:=$ $\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}{\mathbb{P}}(y|g,{V})$
$=$ $\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}\exp\Big(-\|y-\Pi(g\circ{V})\|^{2}/2\sigma^{2}\Big)$
$=$ $\operatorname*{arg\,min}_{g\in\mathsf{SO}(3)}\|y-\Pi(g\circ{V})\|^{2},$ (2.9)
where the first equality follows from (2.5), and the second equality follows from the assumption $\Sigma=\sigma^{2}I_{d\times d}$. In other words, $\hat{g}_{\mathsf{MLE}}$ maximizes the conditional density (2.5) over all possible rotations in $\mathsf{SO}(3)$. In the absence of the tomographic projection (e.g., in cryo-ET), the MLE estimator further simplifies to
$\hat{g}_{\mathsf{MLE}}=\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}y^{\mathsf{T}}(g\circ{V}),$
which corresponds to the rotation $g$ such that the rotated structure maximizes the correlation with $y$. This estimator is also frequently used in single-particle cryo-EM, under the assumption that the norm of $\Pi(g\circ{V})$ is approximately constant for all $g$. This estimator can be approximated by performing a search over a pre-defined grid of 3D rotations, selecting the rotation that minimizes the distance between the measurement $y$ and the projected rotated volume $\Pi(g\circ{V})$. This approach forms the basis of standard practices in single-particle cryo-EM and cryo-ET.
@@ 2.3 The MLE, MAP and MMSE orientation estimators
The MAP estimator extends the MLE estimator, by incorporating prior knowledge on the distribution of the rotation in the special group $\mathsf{SO}(3)$, $\mathrm{d}\Lambda(g)$. Formally, the MAP estimator is defined as
$\hat{g}_{\mathsf{MAP}}$ $=\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}{\mathbb{P}}_{\Lambda}(g|y,{V})$
$=\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}\frac{{\mathbb{P}}(y|g,{V})\mathrm{d}\Lambda(g)}{\int_{\mathsf{G}}{\mathbb{P}}(y|g,{V})\mathrm{d}\Lambda(g)}$
$=\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}{\mathbb{P}}(y|g,{V})\mathrm{d}\Lambda(g)$
$=\operatorname*{arg\,max}_{g\in\mathsf{SO}(3)}\exp\Big(-\|y-\Pi(g\circ{V})\|^{2}/2\sigma^{2}\Big)\mathrm{d}\Lambda(g)$
$=\operatorname*{arg\,min}_{g\in\mathsf{SO}(3)}\|y-\Pi(g\circ{V})\|^{2}/2{\sigma}^{2}-\log\left(\mathrm{d}\Lambda(g)\right),$ (2.10)
where the second equality follows from Bayes’ law (2.4), and the third equality from independence of the denominator on $g$. In other words, $\hat{g}_{\mathsf{MAP}}$ maximizes the posterior density (2.6) over all possible rotations in $\mathsf{SO}(3)$. It can be seen that the MAP estimator $\hat{g}_{\mathsf{MAP}}$ (2.10) coincide with the MLE estimator $\hat{g}_{\mathsf{MLE}}$ (2.9) when the prior distribution $\mathrm{d}\Lambda(g)$ is uniform over $\mathsf{SO}(3)$.
@@
For any rotation distribution $\Lambda$ and loss function $\mathcal{L}(\cdot,\cdot)$ over $\mathsf{SO}(3)$, the MLE and MAP estimators can be further improved by the corresponding Bayes estimator which minimizes the posterior expected loss $\mathbb{E}_{\Lambda}[\mathcal{L}(g,\cdot)\mid y,{V}]$. In particular, following the definition (2.3) of the Bayes estimator, and for the chordal distance $d_{F}(\cdot,\cdot)$ (2.7), with its corresponding squared loss $\mathcal{L}_{F}(\cdot,\cdot)=d_{F}^{2}(\cdot,\cdot)$, the Bayes estimator takes the form
$\hat{g}_{\mathsf{MMSE}}:=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\ \mathbb{E}_{\Lambda}\!\left[d_{F}^{2}\left(g,\hat{g}\right)\mid y,{V}\right],$ (2.11)
where the expectation is taken over the posterior density (2.6). Here we denote the estimator as $\hat{g}_{\mathsf{MMSE}}$, since minimizing the squared chordal distance is equivalent to minimizing the entrywise mean-square error between the two $3\times 3$ rotation matrices. Alternative distance measures, such as the geodesic distance on $\mathsf{SO}(3)$ discussed earlier, could certainly be considered. In this work, however, we focus on the chordal distance, primarily because it admits a particularly convenient derivation of the associated Bayes estimator, as shown below, thereby offering both computational efficiency and interpretability. Importantly, these Bayes estimators are generally distinct from both the MLE and MAP estimators—even under a uniform prior—since it minimizes the posterior expected loss rather than selecting the most probable rotation.
To compute $\hat{g}_{\mathsf{MMSE}}$, we begin by noting that it is defined over the manifold $\mathsf{SO}(3)$. In general, directly solving the optimization problem on a non-Euclidean manifold is challenging. To address this, we first relax the optimization domain to the ambient Euclidean space $\mathbb{R}^{3\times 3}$, which leads to the following intermediate estimator:
$\hat{g}_{\mathsf{relax}}$ $:=\underset{\hat{g}\in\mathbb{R}^{3\times 3}}{\operatorname*{arg\,min}}\ \mathbb{E}_{\Lambda}\!\left[d_{F}^{2}\left(g,\hat{g}\right)\,\middle|\,y,{V}\right]$
$=\mathbb{E}_{\Lambda}[g\mid y,{V}]$
$=\frac{\int_{\mathsf{SO}(3)}g\exp\!\Big(-\tfrac{1}{2}(y-\Pi(g\circ{V}))^{\mathsf{T}}\Sigma^{-1}(y-\Pi(g\circ{V}))\Big)\,\mathrm{d}\Lambda(g)}{\int_{\mathsf{SO}(3)}\exp\!\Big(-\tfrac{1}{2}(y-\Pi(g\circ{V}))^{\mathsf{T}}\Sigma^{-1}(y-\Pi(g\circ{V}))\Big)\,\mathrm{d}\Lambda(g)}$
$=\frac{\int_{\mathsf{SO}(3)}g\exp\!\Big(-\tfrac{1}{2\sigma^{2}}\big\|y-\Pi(g\circ{V})\big\|_{2}^{2}\Big)\,\mathrm{d}\Lambda(g)}{\int_{\mathsf{SO}(3)}\exp\!\Big(-\tfrac{1}{2\sigma^{2}}\big\|y-\Pi(g\circ{V})\big\|_{2}^{2}\Big)\,\mathrm{d}\Lambda(g)}.$ (2.12)
Here, since we naturally represent $g\in\mathsf{SO}(3)$ as a $3\times 3$ rotation matrix, the expectation (or integral) can be interpreted directly in this matrix representation, yielding a $3\times 3$ matrix as the posterior mean matrix [^2] More generally, if one works with different representations of $\mathsf{SO}(3)$, or interprets the expectation as acting on operators (as in Proposition 4.1), the resulting estimator may take different forms.. However, $\hat{g}_{\mathsf{relax}}$ does not, in general, lie in $\mathsf{SO}(3)$ and hence may not be a valid rotation operator. To obtain a feasible solution on the manifold, we subsequently apply the orthogonal Procrustes procedure to project $\hat{g}_{\mathsf{relax}}$ back onto $\mathsf{SO}(3)$, as discussed next.
Given a matrix $A$, to obtain a valid rotation matrix we seek $\Omega\in\mathsf{SO}(3)$ that is closest to $A$ in the Frobenius norm:
@@
This optimization problem is known as the Orthogonal Procrustes problem in the linear algebra literature. It can be solved efficiently using the singular value decomposition (SVD) of $A$ gower2004procrustes .
Thus, applying the Orthogonal Procrustes procedure to $\hat{g}_{\mathsf{relax}}$ projects it back onto $\mathsf{SO}(3)$, yielding a valid rotation matrix. The following proposition establishes that this valid rotation matrix coincides with the MMSE estimator $\hat{g}_{\mathsf{MMSE}}$, thereby completing the procedure for obtaining the MMSE estimate. A detailed proof is provided in Appendix C.